If the probability of passing a road test on the first try is 62%, what is the mean, variance and standard deviation if there are 54 people taking the road test?
Answer by Boreal(15235)

Binomial (2 outcomes)
Mean=np=0.62*54=33.48.
variance=np(1-p)=0.62*54*0.38=12.72
SD=sqrt(12.72)=3.57 (unrounded from above)